I need to find the solution for h. the problem is 2h^2 - h - 3 = 0
Answer by stanbon(26297)

I need to find the solution for h. the problem is
2h^2 - h - 3 = 0
--------------
2h^2 - 3h + 2h -3 = 0
h(2h-3) + (2h-3) = 0
Factor:
(2h-3)(h+1) = 0
h = 3/2 or h = -1
========================
Cheers,
Stan H.